# A Gift from the Integration of Discriminative and Diffusion-based Generative Learning: Boundary Refinement Remote Sensing Semantic Segmentation

### v1

@@ V-C Speeding Up Representation Learning through Incipient Alignment Strategies
Fig. 11 The impact of representation alignment mechanism and alignment strength ($\lambda$) on model performance using CHN6-CUG and Potsdam datasets, divided into four experimental groups. The w/o Align ($\lambda=0$) group represents training without representation alignment; the break control group stops alignment at step 200.
As shown in Fig. 11, mIoU comparisons between non-aligned (w/o Align) and aligned ($\lambda=0.25,0.5$) conditions demonstrate the efficacy of representation alignment. It is evident that introducing the representation alignment strategy significantly accelerates the model’s performance improvement rate in the early stages of training on both datasets.
On the CHN6-CUG dataset, introducing the representation alignment strategy significantly accelerated the model’s performance improvement within the first 200 steps. In contrast, after 1000 steps, the mIoU metric of the non-aligned group gradually surpassed that of the aligned group, indicating that overly strong alignment may inhibit the model’s autonomous ability to model task-specific semantics.
On the Potsdam dataset, representation alignment with $\lambda=0.5$ demonstrated optimal performance across all training stages. In the experiment with $\lambda=0.5$, after stopping the representation alignment, the model showed notably higher convergence efficiency compared to the group without alignment.
In both tasks, all experiments using representation alignment showed a trend of diminishing performance gains or even decline after 400 steps, whereas models without alignment remained more stable. Therefore, we only applied representation alignment in the early training stages to avoid interference with the model’s learning of task-specific semantics during later stages.
V-D Effects of Boundary Tolerance Thresholds in WF-measure
To assess boundary threshold sensitivity in WF-measure, we evaluated three tolerance margins: 1px, 3px, and 5px. The boundary metric WFm was calculated using the SegFormer model on both the Potsdam and Vaihingen datasets, with results shown below:
TABLE IX: Boundary accuracy results (%) of our framework under different boundary thresholds on the Potsdam and Vaihingen datasets
@@ V-D Effects of Boundary Tolerance Thresholds in WF-measure
The results (shown in Table IX) demonstrate that as the threshold increases, the WFm exhibits an overall upward trend, which aligns with the expectation that higher tolerance relaxes boundary matching requirements. Meanwhile, IDGBR maintains steady WFm metric gains across all thresholds, with relatively small variations in improvement magnitude at different threshold levels (2.48%–2.88% for Potsdam and 1.12%–1.28% for Vaihingen). This indicates that the IDGBR framework is insensitive to boundary tolerance changes and demonstrates robust accuracy consistency.
@@ VI Conclusion
In this work, we address the issue of inaccurate boundary delineation in remote sensing semantic segmentation. First, we observe that discriminative learning excels at capturing low-frequency features but exhibits inherent limitations in learning high-frequency features for semantic segmentation. Subsequently, we validate through qualitative experimental analysis and quantitative theoretical proof that the diffusion denoising process significantly enhances the model’s capability to learn high-frequency features. We further observe that these models lack sufficient semantic inference capabilities for low-frequency features when solely guided by the original image. Consequently, we integrate the strengths of discriminative learning and diffusion-based generative learning to propose IDGBR. This framework employs a conditional guidance network to learn joint guidance representation from the coarse segmentation map and the input image, thereby leveraging the respective advantages of discriminative and generative models in semantic and boundary aspects. We conducted extensive experiments on CNN-based architectures, Transformer-based architectures, and self-supervised models in both binary and multi-class remote sensing semantic segmentation scenarios. The results demonstrate that our proposed IDGBR framework exhibits stable boundary optimization capabilities and can consistently refine boundaries of coarse segmentation results generated by different baseline models. We also discuss the crucial roles of conditional guidance and cubic timestep sampling techniques in integrating the advantages of discriminative and generative approaches. Furthermore, we categorize IDGBR’s performance under different completeness levels of coarse segmentation guidance into three levels and systematically discuss the method’s applicability and limitations. Subsequently, experiments validate the effectiveness of regularization techniques in accelerating convergence. Finally, we investigate the stability of our method’s boundary optimization under different boundary tolerance thresholds.
@@
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grants No. 42171376 and No. 42471419) and the High-Level Talent Introduction Project of Inner Mongolia University (Grant No. 10000-A25206020).

### v2

@@ V-D Speeding Up Representation Learning through Incipient Alignment Strategies
To validate the effectiveness of our hyperparameter selection, we analyzed model convergence behaviors under varying alignment strengths ($\lambda\in\{0.25,0.5,0.75\}$). As illustrated in Fig. 10, introducing representation alignment significantly accelerates convergence in the early training stages. However, distinct behavioral patterns emerge as training progresses: excessive regularization (e.g., $\lambda=0.75$) leads to severe performance oscillation and degradation, indicating that overly strong external constraints inhibit the model’s intrinsic capacity to capture task-specific semantics. Furthermore, continuous alignment generally yields diminishing returns or even substantial performance declines compared to the non-aligned baseline in later stages (after 400 steps). Consequently, we identified $\lambda=0.5$ combined with an early stopping strategy as the optimal configuration. These results confirm that the alignment mechanism functions best as initial guidance to leverage prior knowledge for accelerated convergence.
V-E Effects of Boundary Tolerance Thresholds in WF-measure
TABLE XV: Boundary accuracy results of our framework under different boundary thresholds on the Potsdam and Vaihingen datasets
@@ V-E Effects of Boundary Tolerance Thresholds in WF-measure
To evaluate the impact of boundary thresholds in the WF-measure metric, boundary thresholds were set at three different values: 1 pixel (denoted as 1px), 3 pixels (denoted as 3px), and 5 pixels (denoted as 5px). The boundary accuracy of the SegFormer model was calculated on both the Potsdam and Vaihingen datasets, with results shown below:
The results (shown in Table XV) demonstrate that as the threshold increases, the WFm metric exhibits an overall upward trend, which aligns with the expectation that higher tolerance relaxes boundary matching requirements. Meanwhile, IDGBR maintains stable accuracy gains across all thresholds, with relatively small variations in improvement magnitude at different threshold levels (approximately 2.5%–2.9% for Potsdam and 1.1%–1.3% for Vaihingen). This indicates that IDGBR framework is insensitive to changes in boundary tolerance and demonstrates consistent boundary accuracy advantages under various boundary threshold conditions, exhibiting robust boundary optimization capabilities.
@@ VI Conclusion
In this work, we address the issue of inaccurate boundary delineation in remote sensing semantic segmentation. First, we observe that discriminative learning excels at capturing low-frequency features but exhibits inherent limitations in learning high-frequency features for semantic segmentation. Subsequently, we validate through qualitative experimental analysis and quantitative theoretical proof that the diffusion denoising process significantly enhances the model’s capability to learn high-frequency features. We further observe that these models lack sufficient semantic inference capabilities for low-frequency features when solely guided by the original image. Consequently, we integrate the strengths of discriminative learning and diffusion-based generative learning to propose IDGBR. This framework employs a conditional guidance network to learn joint guidance representation from the coarse segmentation map and the input image, thereby leveraging the respective advantages of discriminative and generative models in semantic and boundary aspects. We conducted extensive experiments on CNN-based architectures, Transformer-based architectures, and self-supervised models, alongside comparative benchmarks against boundary optimization methods, covering both binary and multi-class remote sensing semantic segmentation scenarios. The results demonstrate IDGBR’s robust boundary refinement and superiority over boundary optimization methods. We also discuss the crucial roles of conditional guidance and cubic timestep sampling techniques in integrating the advantages of discriminative and generative approaches. Furthermore, we categorize IDGBR’s performance into three levels based on the proposed repair efficacy metric, systematically discussing the method’s applicability and limitations. Subsequently, experiments analyze the effectiveness of regularization techniques and boundary threshold effects in WF-measure.
@@
Supplementary Materials
I. Equation Derivation Details
-A Proof of the Wiener Filter Solution in the Diffusion Probabilistic Model
Here, we present the general mathematical derivation for treating the diffusion model as a Wiener filter. Assume that $z_{0}$ is a wide-sense stationary signal and $\epsilon$ is white Gaussian noise with variance $\sigma^{2}=1$. In the Diffusion Probabilistic Model (DPM), the signal and noise are mixed at any arbitrary timestep $t$, and the resulting observation can be expressed as: $z_{t}=\sqrt{\bar{\alpha}_{t}}\,z_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$ To further simplify the model and obtain an explicit closed-form solution, we introduce a linear denoising Wiener filter $h_{t}$. The objective function for the diffusion model can then be redefined as
$J_{t}=\left\|\sqrt{\bar{\alpha}_{t}}\,z_{0}-h_{t}*z_{t}\right\|^{2}$ (A1)
where $\bar{\alpha}_{t}$ denotes the scaling factor for the signal component at timestep $t$, which decreases as $t$ increases; the symbol “*” denotes the convolution operator.
To facilitate analysis and derive the optimal solution for the filter, we convert the above time-domain convolution operation into the frequency domain. According to the convolution theorem, the time-domain convolution $h_{t}*z_{t}$ is equivalent to the product in the frequency domain, i.e., $\mathcal{F}[h_{t}*z_{t}]=\mathcal{F}[h_{t}]\cdot\mathcal{F}[z_{t}]$, where $\mathcal{F}[\cdot]$ denotes the Fourier transform and $\mathcal{H}_{t}$ represents the corresponding filter in the frequency domain. Thus, the transition from the time-domain convolution to frequency-domain multiplication can be expressed as:
$\mathcal{H}_{t}\left(\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathcal{F}[\epsilon]\right)=\\ \mathcal{F}\left[h_{t}*\left(\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\right)\right]$ (A2)
Furthermore, the error signal $J_{t}(f)$ can be written as
$\mathcal{J}_{t}(f)=\sqrt{\bar{\alpha}_{t}}\,\mathcal{Z}_{0}-\mathcal{H}_{t}\left(\sqrt{\bar{\alpha}_{t}}\,\mathcal{Z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\mathcal{F}[\epsilon]\right)$ (A3)
We obtain the filter by minimizing the mean squared error objective:
$\min_{h}\mathbb{E}\left[|\mathcal{J}_{t}(f)|_{2}^{2}\right]=\min_{h}\mathbb{E}\left[\mathcal{J}_{t}(f)^{*}\mathcal{J}_{t}(f)\right]\\ =\min_{h}\mathbb{E}\left[\left(\sqrt{\bar{\alpha}_{t}}\mathcal{Z}_{0}-\mathcal{H}_{t}\left(\sqrt{\bar{\alpha}_{t}}\mathcal{Z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathcal{F}[\epsilon]\right)\right)\cdot\right.\\ \left.\left(\sqrt{\bar{\alpha}_{t}}\mathcal{Z}_{0}-\mathcal{H}_{t}\left(\sqrt{\bar{\alpha}_{t}}\mathcal{Z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathcal{F}[\epsilon]\right)\right)\right]$ (A4)
where the symbol denotes the complex conjugate. To obtain the Wiener filter that minimizes the mean squared error, we take the derivative of the objective function with respect to $\mathcal{H}_{t}(f)$ and set it to zero. Since $\epsilon$ and $z_{0}$ are conditionally independent, we have:
$\frac{\partial\mathbb{E}[|\mathcal{J}_{t}(f)|_{2}^{2}]}{\partial\mathcal{H}_{t}(f)}=2\mathcal{H}_{t}\left(\bar{\alpha}_{t}|\mathcal{Z}_{0}(f)|^{2}+(1-\bar{\alpha}_{t})|\mathcal{F}[\epsilon]|^{2}\right)\\ -2\bar{\alpha}_{t}|\mathcal{Z}_{0}(f)|^{2}=0$ (A5)
where $|\mathcal{Z}_{0}(f)|^{2}=\mathbb{E}[\mathcal{Z}_{0}\mathcal{Z}_{0}^{*}]$, $|\mathcal{F}[\epsilon]|^{2}=\mathbb{E}[\mathcal{F}[\epsilon]\mathcal{F}[\epsilon]^{*}]$. Since the variance of the white noise is $\sigma^{2}$, $|\mathcal{F}[\epsilon]|^{2}$ is a constant $\sigma^{2}$. Simplifying further, we obtain:
$\mathcal{H}_{t}^{*}(f)=\frac{\bar{\alpha}_{t}|\mathcal{Z}_{0}(f)|^{2}}{\bar{\alpha}_{t}|\mathcal{Z}_{0}(f)|^{2}+(1-\bar{\alpha}_{t})\sigma^{2}}$ (A6)
Here, $|\mathcal{Z}_{0}(f)|^{2}$ denotes the power spectrum of $z_{0}$, and $\mathcal{H}_{t}(f)$ denotes the frequency response of $h_{t}$. According to the commonly assumed power-law distribution of natural image spectra [51], we have $\mathbb{E}[|\mathcal{Z}_{0}(f)|^{2}]\approx 1/f^{2}$. Thus, the Wiener filter in the frequency domain can be expressed as:
$\mathcal{H}_{t}^{*}(f)\approx\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t}+(1-\bar{\alpha}_{t})f^{2}}$ (A7)
II. Network Architecture and Dimensional Specifications
Figure 11: Overview of the IDGBR framework. (a) Data Flow: Macroscopically demonstrates the two-stage inference process from the first-stage discriminative coarse segmentation to the second-stage generative fine optimization. (b) Denoising Unet: Illustrates the hierarchical architecture of the backbone network. (c) Conditional Guidance (CG) Net: Highlights the synergistic mechanism between frozen parameters and trainable modules, as well as the feature injection process. (d) Conditional Guidance (CG) Module: Shows the details of the multi-head cross-attention and zero-convolution structures within the module.
Building upon the schematic diagrams provided in the main text, Fig. 11 further details the dimensional specifications and data flow of the proposed IDGBR framework.
Specifically, panel (a) illustrates the collaborative two-stage inference workflow: Stage I utilizes the trained discriminative model to generate a rough label from the original image; in Stage II, the rough label is first fed into the label encoder ($E_{e}$) to be transformed into continuous features, then concatenated with the original image and compressed by the VAE encoder into latent features to guide the Denoising Unet in generating refined semantic representations, which are ultimately reconstructed into the final high-precision label via the VAE decoder and label decoder ($D_{e}$). Here, $E_{e}$ and $D_{e}$ are introduced to resolve the incompatibility between discrete labels and the continuous latent space. Specifically, $E_{e}$ maps discrete labels to a 3-channel continuous space via an embedding layer, normalized and scaled to the range [1,1][-1,1] using a Sigmoid function; meanwhile, to enhance training robustness, Gaussian noise with a variance of 0.25 is injected into the features during training before they are fed into the auxiliary decoder $D_{e}$ (which maps features back to a $K$-channel probability space constrained by cross-entropy loss).
Panels (b) and (c) illustrate in detail the hierarchical encoding processes within the Denoising Unet and the Conditional Guidance (CG) Net. Both networks progressively downsample feature maps from a resolution of $64\times 64$ to $8\times 8$. Notably, the encoder of the CG Net serves as a pseudo-Siamese branch of the Denoising Unet encoder, parallelly receiving and encoding the dual conditional inputs $\{c_{l},c_{r}\}$. The extracted multi-scale conditional features are then processed by the CG Module and re-injected into the middle layer and corresponding decoder blocks of the Denoising Unet.
Finally, panel (d) reveals the internal details of the Conditional Guidance (CG) Module. Here, the noisy features $f_{\text{noise}}$ are first concatenated with the conditions $\{c_{l},c_{r}\}$ along the channel dimension to form a hybrid feature map of $3C\times H\times W$, which is processed by a zero-convolution layer to serve as the Query; simultaneously, the dual conditions are concatenated into a $2C\times H\times W$ feature vector to act as the Key and Value for the multi-head cross-attention mechanism. Ultimately, the attention output is fused with the query via a residual connection.
III. Experimental Details
-A Qualitative Results on the FGFD.
FGFD is primarily used to validate the boundary optimization capability of the IDGBR framework in complex texture scenes.
Figure 12: Qualitative results of different methods on FGFD.
The complex spatiotemporal attributes of farmland pose significant challenges to traditional discriminative models for segmentation. As shown in Fig. 12, the visualization results collectively demonstrate the distinguishability of farmland and the continuity of field roads in three farmland scenarios. Fig. 12 (a) shows typical northern dryland farmland in Shanxi. For non-cultivated areas that are difficult to distinguish, even when patches appear regular in the original imagery, discriminative models produce patches with incomplete geometric structures and fragmented boundaries. Our method effectively suppresses low-confidence artifact regions, significantly improving patch boundary integrity and standardization. Fig. 12 (b) depicts highland grassland in Tibet. In the original imagery, shadows cast by trees along farmland boundaries create light-dark divisions within the plots. Discriminative models are prone to boundary misclassification errors in such areas. Our method mitigates segmentation errors of baseline models in shadow regions, enhancing the accuracy and continuity of farmland boundary recognition. Fig. 12 (c) showcases composite farmland in Jiangxi, where farmland boundaries are often obscured by dense vegetation cover, resulting in blurred boundaries. Due to the minimal spectral differences between vegetation and farmland in the original image, discriminative models struggle to accurately delineate boundary positions, leading to boundary ambiguity and discontinuous segmentation. Our method corrects boundary blurring caused by vegetation interference, improving the geometric consistency and completeness of segmentation results in complex regions.
-B Qualitative Results on the WHU Building Dataset
To validate the effectiveness of IDGBR in building boundary extraction and cross-temporal generalization, we conducted experiments on the bi-temporal images of the WHU Building Change Detection dataset.
Figure 13: Qualitative analysis on the WHU Building post-disaster phase dataset.
Figure 14: Qualitative analysis on the WHU Building post-reconstruction phase dataset.
Figs. 13 and 14 demonstrate the qualitative optimization results of IDGBR on post-disaster and post-reconstruction phase images from the WHU Building dataset, respectively. It can be observed that the original discriminative models generally exhibit significant errors near building boundaries, including local false responses, irregular boundary contours, and insufficient geometric structural integrity, particularly in areas with complex roof shapes or blurred boundaries. After introducing IDGBR, all models show significantly reduced errors at boundaries, with smoother and more coherent building edges, and well-preserved detailed morphology, demonstrating the method’s clear advantages in boundary modeling. Furthermore, the cross-temporal transfer experiments in Fig. 14 verify that models optimized with IDGBR maintain high boundary accuracy during the reconstruction phase, significantly suppressing segmentation errors caused by cross-temporal differences while maintaining stable boundary morphology. This indicates that IDGBR possesses strong temporal stability and cross-scene adaptation capabilities.
-C Qualitative Results on Multi-Class Segmentation Tasks
To further verify the generalization of our method in complex multi-category scenarios, we visualize the segmentation results on the Potsdam and Vaihingen datasets.
Figure 15: Qualitative results of different methods on multi-class segmentation datasets (Potsdam and Vaihingen).
-D Inference Efficiency Analysis
TABLE XVI: Performance comparison of cubic sampling strategy on binary datasets with different foreground ratios.
Dataset Cubic Strategy IoU F1-score OA Kappa WFm (3px B) Foreground Ratio (%)
CHN6-CUG $\checkmark$ 62.96 77.27 97.44 75.91 37.16 7.59
$\times$ 64.54 78.45 97.50 77.12 37.64
FGFD $\checkmark$ 87.52 93.35 92.32 84.27 49.24 59.34
$\times$ 87.66 93.42 92.39 84.40 49.78
WHU (Post-disaster Phase) $\checkmark$ 94.76 97.31 99.19 96.83 76.46 15.00
$\times$ 94.72 97.29 99.18 96.81 75.91
TABLE XVII: Comparison of GFLOPs and parameters with Conditional Diffusion Model.
Methods GFLOPs Params (M)
Control SD (25 steps) 24576.92 1427.51
Ours (25 steps) 29121.39 1521.09
[…]
Table XVII presents a comparative analysis of the computational complexity of IDGBR against a conditional diffusion baseline. Diffusion-based methods require an iterative denoising process, which inherently leads to substantial computational overhead. Specifically, compared to the standard conditional diffusion baseline, IDGBR introduces a moderate increase of 18.5% in GFLOPs and 6.6% in parameters. The justification for this increased computational cost should be evaluated within the broader context of the full remote sensing mapping lifecycle. This workflow spans from image acquisition and automated interpretation to manual verification, graphic editing, and final map publication. Low-quality segmentation boundaries—manifested as blurring, jagged edges, or topological errors—directly increase the downstream burden of manual verification and editing. By significantly improving boundary quality at a modest computational cost, our method replaces a costly and uncertain manual editing stage with a deterministic, one-time computational investment. Consequently, the overall cost of the mapping process is reduced.
-E Applicability Analysis of Cubic Sampling Strategy in Binary Segmentation Tasks
Experimental results presented in the main text indicate that the cubic sampling strategy resulted in a slight performance degradation on the CHN6-CUG dataset. Given the extremely low proportion of foreground targets in this dataset (approximately 7.59%), we preliminarily hypothesized that this phenomenon stemmed from the severe imbalance in sample distribution. To investigate this, we introduced the FGFD dataset (with a balanced foreground ratio of 59.34%) and the WHU dataset (with a similarly low foreground ratio of 15.00%) for comparative experiments, as detailed in Table XVI.
The experimental results reveal that there is no linear correlation between the foreground ratio and the efficacy of the cubic sampling strategy. On the WHU dataset, which features a similarly sparse foreground, cubic sampling yielded a slight performance improvement. Conversely, on the foreground-dominated FGFD dataset, the strategy exerted a suppressive effect on performance. This comparison demonstrates that the performance variations of cubic sampling across datasets cannot be fully attributed solely to the proportion of ground object classes; its performance is affected by additional factors that are not yet fully understood. In binary segmentation tasks, it should be regarded as a condition-sensitive optional optimization strategy.